OpenStudy (anonymous):

The acceleration function (m/s^2) and the initial velocity are given for a particle moving along a line. Find each of the following: a(t)=2t+3, v(0)=-4, 0 is less than or equal to t which is less than or equal to 3 A)the velocity at time t B) The distance traveled during the given time interval. C) Show that for motion in a straight line with constant acceleration a, and initial velocity v0, and initial displacement s0, the displacement after time t is S=1/2at^2+v0t+s0 Im confused what to do for C

OpenStudy (michele_laino):

hint: in order to find the space traveled, at time t, we have to integrate twice the acceleration a(t)

OpenStudy (michele_laino):

whereas in order to find the velocity, we have to integrate once the function a(t)

OpenStudy (michele_laino):

for example, the velocity v(t), is given by the subsequent computation: \[\Large v\left( t \right) = \int {a\left( t \right)dt = } \int {\left( {2t + 3} \right)dt} = ...?\]

OpenStudy (anonymous):

v(t)=t^2+3t-4 and then s(t)=t^3/3 +3/2t-4t

OpenStudy (michele_laino):

v(t) is right, whereas integrating once v(t) I get: \[\Large s\left( t \right) = \frac{{{t^3}}}{3} + \frac{3}{2}{t^2} - 4t + k\] where k is an arbitrary constant, whose meaning is the space at t=0, namely: k=s(0)=s_0